OpenStudy (anonymous):

Mary had 90 fliers to post around town. Last week, she posted 1/6 of them. This week, she posted 1/5 of the remaining fliers. How many fliers has she still not posted?

OpenStudy (ajprincess):

what is 1/6 of 90?

OpenStudy (anonymous):

i have no clue, im horrible at this stuff

OpenStudy (ajprincess):

\(\frac{1}{6}*90=\large\frac{90}{6}=?\)

OpenStudy (anonymous):

15

OpenStudy (anonymous):

the answer is 60?

OpenStudy (ajprincess):

ya

OpenStudy (ajprincess):

The number of fliers posted last week=15. Remaining fliers=90-15=75 Number of fliers posted this week=1/5 of 75=15 Remaining fliers=75-15=60

OpenStudy (anonymous):

1/6 of 90= 90/6=15 so the remaining is 90-15=75 1/5 of 75=75/5=15

OpenStudy (anonymous):

and 75-15 = 60